EDITORIAL

Terry Goodman

We at MJMS hope your year has started well. We appreciate your continued interest in and support of our journal and encourage you to share it with colleagues. We also encourage you to share your mathematical ideas by submitting potential short articles and problems such as those that are listed on the inside cover of each issue of MJMS. In particular, I am consistently amazed at the continuing developments in technology and the implications these developments have for the very nature of mathematics itself, how we learn/teach mathematics, and how we conduct mathematics research.

I recently had the opportunity to visit with several teachers and observe some exciting classes they were developing. In one instance, a calculus course and a physics course have been combined into a two hour block course. The teachers team teach this block course with an emphasis on student–centered laboratory activities and projects. The students often work in small groups using computers as their primary investigative tool.

The students in this course make use of graphing, simulation, statistical, and powerful mathematics software as they investigate a variety of concepts and applications. On some days the focus in the course is on mathematics, some days on physics, but most often the students are using calculus concepts and techniques to explore applications in physics. Other students are working on a variety of other mathematics and science projects. Internet is used to help the students communicate with other students around the country as well as obtain information from mathematicians and scientists.

Other courses are in the process of being developed and they too will emphasize student– centered curricula, investigative laboratory techniques, and the use of technology to implement these investigations.

We at Central Missouri State University are considering ways that we might collaborate with other schools to further develop these and similar programs. Specifically, we want to find ways to involve our students in working with such programs. We are also considering ways that we can further incorporate the use of technology in our programs on campus. There are some real possibilities in this area!